MONT 104N – Modeling the Environment Second Essay Assignment Due: No later than 5:00pm on Friday, November 16 November 6, 2018

The second essay assignment for the course will be based on another short essay called Hidden Intellectualism starting on p. 248 of They Say, I Say (4th edition). You should start by reading this carefully and thinking about what Gerald Graff is saying about the ways conventional education can miss opportunities for reaching students who don’t always respond to school work but who might be passionately and thoughtfully engaged with other things.

This assignment will be more of an opportunity to reflect on what the author is saying and relate that to your own experiences up to this point in your education (and maybe especially your experiences so far in college). As you will see, it’s not an assignment related to the mathematical or scientific questions we have been considering in class.

You should structure your essay according to the “They Say, I Say” paradigm. In other words, one possible structure would be to start by laying out in detail what Graff is saying and then transition to your reaction and how you think your own experiences either do or do not correspond to his main point(s). Another possible approach would be to go through the main ideas from his essay one by one and respond to each of them in turn. Both ways would be possible; it’s up to you to find what works for what you want to say.

Feel free to make this one as personal and informal or colloquial as you want. I want to hear what you have to say about this, in your own voice. Please don’t make this into some dry, academic, content-free, page-filling exercise. For that reason, there is no fixed page quota for your essay. Write as much or as little as you want. The only requirement is that you give a good summary of Graff’s ideas along the way and respond to them in a meaningful way.

Here are some questions to get you started if you don’t know what to say: First, exactly what does Graff mean by “intellectualism?” What constitutes the quality of being an “intellectual” for him? (Please don’t use a dictionary definition in your essay!) Is being an “intellectual” anything you have ever consciously considered for yourself?

If you are a varsity athlete, do you see a connection between what Graff is claiming about participation in sports or other “non-intellectual” activities and what you get out of what you do? Or is he “all wet” about the way sports and other pursuits can be “intellectualism by other means” (p. 250). Would anyone who is not an intellectual already think to say that? If you are not an athlete and you want to argue that traditional strictly academic work (perhaps reading the “classics”) is more valuable than any focus on elements of popular culture or sports, by all means go ahead and do that too.

Your sentences should be clear, readable, and not overly wordy. You should aim for a lively style and the writing should be in your own voice. But please think twice before using very colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

Errors in mechanics (spelling, punctuation, grammar) and usage should not annoy the reader (i.e. Prof. Little). Try as much as possible to avoid those mistakes.